3x + 2y = -10
x= 1 - 5y
SUBTRACTION METHOD

Answer by stanbon(75887)

3x + 2y = -10
x= 1 - 5y
-----
Rearrange:
3x + 2y = -10
x + 5y = 1
------
Modify for subtraction:
3x + 2y = -10
3x + 15y= 3
------
Subtract and solve for "y":
13y = -13
y = -1
-----
Solve for "x":
x = 1 - 5y
x = 1 - 5*-1
x = 1 + 5
x = 6
------------------
Cheers,
Stan H.
==================